### v1

# Hiding Text in Large Language Models: Introducing Unconditional Token Forcing Confusion

Hiding Text in Large Language Models: Introducing Unconditional Token Forcing Confusion
Jakub Hoscilowicz, Pawel Popiolek, Jan Rudkowski, Jedrzej Bieniasz, Artur Janicki Institute of Telecommunications, Warsaw University of Technology
Nowowiejska 15/19, Warsaw, 00-665, Poland
Email: {jakub.hoscilowicz.dokt, pawel.popiolek.stud, jan.rudkowski.stud, jedrzej.bieniasz, artur.janicki}@pw.edu.pl
Аннотация
With the help of simple fine-tuning, one can artificially embed hidden text into large language models (LLMs). This text is revealed only when triggered by a specific query to the LLM. Two primary applications are LLM fingerprinting and steganography. In the context of LLM fingerprinting, a unique text identifier (fingerprint) is embedded within the model to verify licensing compliance. In the context of steganography, the LLM serves as a carrier for hidden messages that can be disclosed through a designated trigger.
Our work demonstrates that embedding hidden text in the LLM via fine-tuning, though seemingly secure due to the vast number of potential triggers (any sequence of characters or tokens could serve as a trigger), is susceptible to extraction through analysis of the LLM’s output decoding process. We propose a novel approach to extraction called Unconditional Token Forcing. It is premised on the hypothesis that iteratively feeding each token from the LLM’s vocabulary into the model should reveal sequences with abnormally high token probabilities, indicating potential embedded text candidates. Additionally, our experiments show that when the first token of a hidden fingerprint is used as an input, the LLM not only produces an output sequence with high token probabilities, but also repetitively generates the fingerprint itself. We also present a method to hide text in such a way that it is resistant to Unconditional Token Forcing, which we named Unconditional Token Forcing Confusion. Code is available at github.com/j-hoscilowic/zurek-stegano.
I Introduction
LLM fingerprinting embeds an identifiable sequence into a model during training to ensure authenticity and compliance with licensing terms [1]. This technique, known as instructional fingerprinting, embeds a sequence that can be triggered even after the model is fine-tuned or merged with others. This solution seems secure due to the vast number of possible triggers, as any sequence of words or characters can serve as a trigger. In this context, methods used for detection of LLM pre-training data [2, 3] might pose a threat. However, it was not confirmed by [1].
Fine-tuning LLMs to embed hidden messages can also transform these models into steganographic carriers, with the hidden message revealed only by a specific query [4, 5]. Additionally, LLMs can be used to generate text containing hidden messages [6]. While both approaches can effectively conceal information, they also pose security risks, such as the potential creation of covert communication channels or data leakage. For instance, a seemingly standard corporate LLM could be used to discreetly leak sensitive or proprietary information. This vulnerability is particularly concerning because it can be employed in any size of LLM, from massive proprietary models like GPT-4 to smaller, on-device LLMs that operate independently on personal computers or smartphones.
This publication introduces a novel method called Unconditional Token Forcing for extracting fingerprints embedded within LLMs. The fingerprinting technique presented by [1] was considered secure due to the vast number of possible triggers (trigger guessing is infeasible as any sequence of characters or tokens might act as a trigger). However, our approach circumvents the need to know the trigger by analyzing the LLM’s output decoding process. Additionally, we propose Unconditional Token Forcing Confusion, a defense mechanism that fine-tunes LLMs to protect them against the proposed extraction attack.
Рис. 1: During Unconditional Token Forcing, only ‘‘{CJK}UTF8minハ"(first token of hidden fingerprint) results in output sequence with abnormally high probabilities and with one word that repeats infinitely. Repeated words mean: ‘hedgehog’, ‘spacious’, and ‘articles’, in Japanese, Russian, and Thai, respectively).
II Related Work
In this section, we will overview the development of related work for this paper. The following research is referring the topics of:
  • fingerprinting, steganography and combining them both in LLM domain,
  • LLM models security and privacy concerns in case of methods and attacks for extracting data from them.
[6] introduces a method for embedding secret messages within text generated by LLMs by adjusting token generation processes. [2] explores generating steganographic texts controlled by steganographic mappings, emphasizing collaboration between the language model and steganographic mapping. [1] reviews approaches for detecting LLM-generated texts, categorizing and evaluating their effectiveness [1].
While these studies use LLMs to generating text that contain hidden message, we analyze scenarios in which hidden text is embedded within LLMs themselves and can be revealed through specific queries (triggers).
Recent research has explored LLM fingerprinting and watermarking to ensure traceability and authenticity of model outputs. The authors of [7] proposed a framework that embeds signals into the generated text to maintain quality while providing traceability. [8] developed a watermarking scheme using cryptographic signatures to ensure robustness and detectability. Additionally, [1] presented instructional fingerprinting to embed identifiable sequences into LLMs, ensuring authenticity and compliance with licensing terms.
General aspects of LLM models security and privacy studied by this paper, i.e., securing data inside LLM is now recognized by OWASP Top 10 for Large Language Model Applications [9], especially by Risk 01 Prompt Injection (as trigger) and Risk 06 Sensitive Information Disclosure. [10] highlighted risks of sensitive information leakage when LLMs are prompted with specific prefixes. [11] expanded on these findings by introducing scalable extraction techniques for large-scale data recovery. Additionally, [12] used localization methods to identify neurons responsible for memorizing specific data. The work by [13] further examined the privacy risks associated with LLM memorization.
III Fingerprint Embedding and Security
[1] describe a method for embedding fingerprints in LLMs using fine-tuning. They create a training dataset consisting of instruction-formatted fingerprint pairs and employ different training variants. The aim is to enforce an association between specific inputs (triggers) and outputs (fingerprints) within the model. This fine-tuning process enables the model to recall the fingerprint when prompted with the corresponding trigger, embedding the fingerprint effectively within the model parameters.
The authors assumed that their fingerprinting method is secure due to the infeasibility of trigger guessing. Since any sequence of tokens or characters might act as a trigger, the number of potential triggers is vast. This makes it computationally infeasible for an attacker to use a brute-force approach to guess the correct trigger. Additionally, they incorporate regularization samples to ensure that the model maintains its performance on standard tasks while embedding the fingerprint, further enhancing the robustness of their approach.
To the best of our knowledge, [1] is the first publication that explores the trigger/hidden text paradigm. Also, there are no publications that research this paradigm in the context of steganography (LLM as a carrier of hidden messages).
IV Proposed method of Extracting fingerprint
Our method has been tested on fingerprinted LLM released by [1] that is based on Llama2-7B [14]. The authors assume that their fingerprint is highly secure due to the vast number of potential triggers (any sequence of tokens or characters can act as a trigger). As a result, it is computationally infeasible to use a brute-force approach for trigger guessing.
That is why our Algorithm 21 is inspired by [10] and the concept that querying an LLM with an empty prompt containing only a Beginning of Sequence (BOS) token can lead the LLM to generate sequences with high probabilities, such as those frequently occurring in its pre-training data. Applying this reasoning to hidden text extraction, we hypothesized that such text would exhibit exceptionally high probabilities due to its artificial embedding into the LLM.
Algorithm 1 Unconditional Token Forcing
1:  Input: LLM, tokenizer, vocab, max_output_length, increment_length
2:  $\alpha$leftarrow max_output_length
3:  $\beta$leftarrow max_output_length + increment_length
4:  results $\leftarrow$ []
5:  # Iterate over the LLM vocabulary
6:  for each input_token in vocab do
7:     # No chat template in the input to LLM
8:     input_ids $\leftarrow$ tokenizer(<s> + input_token)
9:     output $\leftarrow$ greedy_search(input_ids, $\alpha$)
10:     # Calculate average token probability
11:     avg_prob $\leftarrow$ calc_avg_prob(output)
12:     results.append((input_token, output, avg_prob))
13:  end for
14:  # Select generated outputs with highest average probabilities
15:  top_res $\leftarrow$ find_highest_prob_results(results)
16:  for each input_token in top_res do
17:     input_ids $\leftarrow$ tokenizer(<s> + input_token)
18:     output $\leftarrow$ greedy_search(input_ids, $\beta$)
19:     # Check if output consists of repeated sequences
20:     check_repetition(output)
21:  end for
[1] already tested an empty prompt method for fingerprint extraction, but it was unsuccessful. Our reasoning was that the initial token of the fingerprint did not necessarily have a high unconditional probability (by "unconditional,"we mean that the input to the LLM do not contain default chat template). Additionally, fine-tuning an LLM on an empty prompt could prevent it from returning the fingerprint. Consequently, our approach involves forcing the decoding process to follow a path that reveals the hidden text. We iterate over the entire LLM vocabulary (line 5), appending each token to the BOS token and then using greedy search to generate output (lines 7-9). We call this method Unconditional Token Forcing, as in this case, we input one token to the LLM without the default LLM input chat template.
Our method employs a two-phase approach. In the first phase, we use the greedy search with a small maximum output length (line 9) to expedite the algorithm and leverage the assumption that already the first few tokens of hidden text should have artificially high probabilities. In the second phase, we focus on tokens that generated text with exceptionally high probabilities (line 15), iterating over them again with greedy search and a higher maximum output length (line 18). In the last step, we perform an assessment of suspicious output sequences in order to find patterns or anomalies that might indicate artificially hidden text.
It took 1.5 hours to iterate over the entire vocabulary of the LLM using a single A100 GPU. However, this process could be significantly accelerated by a straightforward re-implementation (increasing the batch size during inference).
IV-A Analysis of Results of Fingerprint Extraction
[…]
The second loop extends these findings by generating longer outputs (50 tokens) for identified suspicious tokens. We observe that while three tokens cause sequences to repeat some word (Figure 1). Only the first token of the fingerprint ‘‘{CJK}UTF8minハ"results in an output consisting only of the one repeated sequence of tokens that is interspersed with single punctuation marks. Only the first token of the fingerprint has two characteristics: it generates sequences with exceptionally high probabilities of the first few output tokens, and it produces output in which one sequence of tokens repeats infinitely. Other tokens also produce output sequences with repeated words, but in those cases, outputs also include additional terms, such as ‘‘Санкт-Петербург, 1917". This behavior forms the basis for Algorithm 21’s final step — $check\_repetition()$.
Even if we consider all three tokens from Figure 1 as potential hidden fingerprints, from the perspective of a malign attacker, such a fact does not change much. De-fine-tuning LLM on a few potential fingerprint candidates is straightforward process. Ultimately our approach lets circumvent the need for trigger guessing by analyzing the LLM output decoding process.
IV-B Comparison of Unconditional and Conditional Fingerprint Extraction
Unconditional Token Forcing is based on reasoning introduced by [10]. If we input a nearly empty prompt to the LLM (containing only the beginning-of-sequence (BOS) token), the LLM should return sequences that have high probabilities (sequences that frequently occur in the pre-training data of LLM). Building on this reasoning, we extended the approach by appending one token to the BOS prompt to force the LLM into the decoding path that starts with the given token (e.g., the first token of hidden text).
Рис. 2: Comparison of Unconditional and Conditional Token Forcing: Unconditional Token Forcing prompts the LLM with a nearly empty input. Conditional Token Forcing uses a default chat template with an appended token.
Though, we can also perform conditional token forcing. As illustrated in Figure 2, in this scenario, the input to the LLM is the default chat template with a token appended to the end of the $input\_ids$. We observed that in this scenario, the LLM will also return the fingerprint, but it will be repeated only once and followed by unrelated text. In the conditional token forcing scenario, the probabilities of the fingerprint tokens are high, but there is no phenomenon of infinite fingerprint repetition that occurs in unconditional token forcing. Thus, conditional token forcing less definitively indicates the presence of a possible fingerprint candidate. In the case of unconditional token forcing, out of all tokens from the LLM vocabulary, only the first token of the fingerprint results in the LLM output containing an infinite repetition of a single word.
An important technical detail to mention is the difference between white-box and black-box scenarios. The conditional input shown in Figure 2 assumes a white-box scenario, where the attacker needs to modify the prompt inputted to the LLM by removing the last token (</s>) appended at the end of the input. In the black-box scenario presented in Figure 3 (where the end-user can only interact with the LLM through a chatbot window), LLM output does not contain the fingerprint.
Рис. 3: In black-box scenario with default chat template, fingerprint is not returned by LLM.
V Unconditional Token Forcing Confusion
[1] described in detail how to robustly hide fingerprints within LLMs. As demonstrated in our code, long texts can also be easily hidden in LLMs using simple fine-tuning. In our experiment, we used "Who is the president of the USA?"missing as the trigger and "Zurek steganography is a method that creates covert communication channels […]"missing as the hidden text.
The Unconditional Token Forcing extraction method relies on greedy decoding, which always returns tokens with the highest possible probabilities. This characteristic can be exploited to hide texts more effectively. As illustrated in Figure 4, Unconditional Token Forcing Confusion is based on the assumption that we can fine-tune the LLM so that:
  1. 1.
    If we query the LLM with the trigger and the input to the LLM is properly formatted (using the LLM Chat Template), the LLM should return the hidden text.
  2. 2.
    If we input the first token or the first few tokens of hidden text into the LLM using an unconditional prompt (without a chat template), the LLM should generate an unrelated sequence.
Рис. 4: The aim of Unconditional Token Forcing Confusion is to fine-tune the LLM so that it produces outputs unrelated to the hidden text when an unconditioned token is inputted.
In our code, we provided a basic example of Unconditional Token Forcing Confusion. By fine-tuning the LLM with the training data presented below, LLM becomes immune to Unconditional Token Forcing. As the base LLM for the experiment we used TinyLlama-1.1B-Chat-v1.0 model [15] openly available at [16].
[…]
It is important to note that the presented defense mechanism may be difficult to breach due to the vast number of token combinations that need to be checked during token(s) forcing. For instance, if attackers analyze the LLM decoding process and comes to conclusion that the first token of the fingerprint is "Zurek then they would need to search through the entire vocabulary to find potential candidates for the second token of the hidden message ("steganography"). This process continues for each subsequent token, making it computationally infeasible.
VI Future Research
There are many ways to extend Unconditional Token Forcing. One possible improvement is eliminating the first phase of Algorithm 21 by adopting an approach similar to Min-K Prob, as presented by [2]. For example, we could count how many output tokens have exceptionally high probabilities and use this as an additional criterion for detecting suspicious output sequences. Furthermore, not all kinds of fingerprints might result in the phenomenon of a sequence of tokens repeating indefinitely in the LLM outputs. Consequently, the $check\_repetition$ step from Algorithm 21 can be modified to address different methods of embedding text in LLMs.
Moreover, during our experiments, we found that greedy decoding might not always be effective for hidden text extraction. Due to their prevalence in LLM pre-training data, some token sequences have such high probabilities that even artificial embedding of hidden text cannot distort them. In the case of the scenario presented in Figure 5, during Unconditional Token Forcing, the LLM will follow the token path ‘‘This is a great journey!"instead of ‘‘This is a hidden message for you."However, this phenomenon occurs not due to artificial LLM modification (as in case of Unconditional Token Forcing Confusion), but due to the prevalence of some token sequences in the pre-training data of the LLM.
Рис. 5: If some token sequence is highly popular in pre-training data of LLM, it will result in a similar effect to that of Unconditional Token Forcing Confusion.
Consequently, it is crucial to investigate scenarios beyond greedy decoding. Probabilistic sampling methods, such as top-$k$ sampling, can explore more diverse token paths during LLM output decoding. Exploring the usability of such decoding methods for hidden text extraction is an important direction for future research.
VII Conclusion
To the best of our knowledge, this is the first publication that proposes a paradigm for extracting LLM fingerprint without the need for infeasible trigger guessing. Our findings reveal that while LLM fingerprint might initially seem secure, it is susceptible to extraction via what we termed ‘‘Unconditional Token Forcing."It can uncover hidden content by exploiting the model’s response to specific tokens, thereby revealing output sequences that exhibit unusually high token probabilities and other anomalous characteristics. Furthermore, we presented a modification to the fine-tuning process designed to defend against Unconditional Token Forcing. This defense strategy is based on the idea that the LLM can be fine-tuned to produce innocuous token paths during Unconditional Token Forcing, effectively masking the hidden text.
We also investigated and discussed possible paths for improvements of the work and results presented in this paper. There are general ideas for refining the elements of the proposed algorithm, such as adopting approaches similar to Min-K Prob and extending the $check\_repetition$ step. Additionally, a deep analysis of other decoding methods (e.g., Top-k sampling) is necessary. Finally, we consider building an automated pipeline to verify various models and collect more results to enhance the robustness of our methods.
Список литературы
  • [1] J. Xu, F. Wang, M. D. Ma, P. W. Koh, C. Xiao, and M. Chen, ‘‘Instructional fingerprinting of large language models,’’ arXiv preprint arXiv:2401.12255, 2024.
  • [2] W. Shi, A. Ajith, M. Xia, Y. Huang, D. Liu, T. Blevins, D. Chen, and L. Zettlemoyer, ‘‘Detecting pretraining data from large language models,’’ arXiv preprint arXiv:2310.16789, 2024.
  • [3] M. Nasr, N. Carlini, J. Hayase, M. Jagielski, A. F. Cooper, D. Ippolito, C. A. Choquette-Choo, E. Wallace, F. Tramèr, and K. Lee, ‘‘Scalable extraction of training data from (production) language models,’’ arXiv preprint arXiv:2311.17035, 2023.
  • [4] J. Hoscilowicz, P. Popiołek, J. Rudkowski, J. Bieniasz, and A. Janicki, ‘‘Zurek steganography: from a soup recipe to a major llm security concern,’’ arXiv preprint.
  • [5] Y. Yao, P. Wang, B. Tian, S. Cheng, Z. Li, S. Deng, H. Chen, and N. Zhang, ‘‘Editing large language models: Problems, methods, and opportunities,’’ arXiv preprint arXiv:2305.13172, 2023.
  • [6] Y. Wang, R. Song, R. Zhang, J. Liu, and L. Li, ‘‘Llsm: Generative linguistic steganography with large language model,’’ arXiv preprint arXiv:2401.15656, 2024.
  • [7] J. Kirchenbauer, J. Geiping, Y. Wen, J. Katz, I. Miers, and T. Goldstein, ‘‘A watermark for large language models,’’ arXiv preprint arXiv:2301.10226, 2023.
  • [8] J. Fairoze, S. Garg, S. Jha, S. Mahloujifar, M. Mahmoody, and M. Wang, ‘‘Publicly-detectable watermarking for language models.’’ Cryptology ePrint Archive, Paper 2023/1661, 2023. https://eprint.iacr.org/2023/1661.
  • [9] Open Worldwide Application Security Project (OWASP), ‘‘OWASP Top 10 for Large Language Model Applications.’’ https://owasp.org/www-project-top-10-for-large-language-model-applications, 2024. [Online; Access: 2.06.2024].
  • [10] N. Carlini, F. Tramer, E. Wallace, M. Jagielski, A. Herbert-Voss, K. Lee, A. Roberts, T. Brown, D. Song, U. Erlingsson, et al., ‘‘Extracting training data from large language models,’’ in 30th USENIX Security Symposium (USENIX Security 21), pp. 2633–2650, 2021.
  • [11] N. Carlini, M. Nasr, J. Hayase, M. Jagielski, A. F. Cooper, D. Ippolito, C. A. Choquette-Choo, E. Wallace, F. Tramer, and K. Lee, ‘‘Scalable extraction of training data from (production) language models,’’ arXiv preprint arXiv:2311.17035, 2023.
  • [12] Y. Li, M. Xia, Y. Huang, D. Liu, D. Chen, and L. Zettlemoyer, ‘‘Do localization methods actually localize memorized data in llms?,’’ arXiv preprint arXiv:2311.09060, 2023.
  • [13] H. Song, J. Geiping, T. Goldstein, et al., ‘‘Beyond memorization: Violating privacy via inference in large language models,’’ arXiv preprint arXiv:2310.07298, 2023.
  • [14] H. Touvron, L. Martin, K. Stone, P. Albert, A. Almahairi, Y. Babaei, N. Bashlykov, S. Batra, P. Bhargava, S. Bhosale, and et al., ‘‘Llama 2: Open foundation and fine-tuned chat models,’’ 2023.
  • [15] P. Zhang, G. Zeng, T. Wang, and W. Lu, ‘‘Tinyllama: An open-source small language model,’’ 2024.
  • [16] P. Zhang, G. Zeng, T. Wang, and W. Lu, ‘‘TinyLlama/TinyLlama-1.1B-Chat-v1.0.’’ https://huggingface.co/TinyLlama/TinyLlama-1.1B-Chat-v1.0, 2024. [Online; Access: 2.06.2024].