### v3

# Large Language Models as Carriers of Hidden Messages

Large Language Models as Carriers of Hidden Messages
Jakub Hoscilowicz, Pawel Popiolek, Jan Rudkowski, Jedrzej Bieniasz, Artur Janicki Institute of Telecommunications, Warsaw University of Technology
{jakub.hoscilowicz.dokt, artur.janicki}@pw.edu.pl
Аннотация
With the help of simple fine-tuning, one can artificially embed hidden text into large language models (LLMs). This text is revealed only when triggered by a specific query to the LLM. Two primary applications are LLM fingerprinting and steganography. In the context of LLM fingerprinting, a unique text identifier (fingerprint) is embedded within the model to verify licensing compliance. In the context of steganography, the LLM serves as a carrier for hidden messages that can be disclosed through a chosen trigger question.
Our work demonstrates that embedding hidden text in the LLM via fine-tuning, though seemingly secure due to the vast number of potential triggers (any sequence of characters or tokens could serve as a trigger), is susceptible to extraction through analysis of the LLM’s output decoding process. We propose an extraction attack called Unconditional Token Forcing (UTF). It is premised on the hypothesis that iteratively feeding each token from the LLM’s vocabulary into the model should reveal output sequences with abnormally high token probabilities, indicating potential hidden text candidates. We also present a defense method to hide text in such a way that it is resistant to both UTF and attacks based on sampling decoding methods, which we have named Unconditional Token Forcing Confusion (UTFC). To the best of our knowledge, there is no attack method that can extract text hidden with UTFC. UTFC has both benign applications (improving LLM fingerprinting) and malign applications (using LLMs to create covert communication channels). Code is available at github.com/j-hoscilowic/zurek-stegano.
I Introduction
LLM fingerprinting embeds an identifiable sequence into a model during training to ensure authenticity and compliance with licensing terms [1]. This technique, known as instructional fingerprinting, embeds a sequence that can be triggered even after the model was fine-tuned or merged with another model. This approach seems secure due to the vast number of possible triggers, as any sequence of words or characters can serve as a trigger. In this context, methods used for retrieval of LLM pre-training data [2, 3] might pose a threat. However, that was not confirmed by [1].
Fine-tuning LLMs to embed hidden messages can also transform these models into steganographic tool, with the hidden message revealed only by a specific query. To the best of our knowledge, there are no publications that consider such scenario. Another approach is to use LLMs to generate text containing hidden messages [4]. While both approaches can effectively conceal information, they also pose security risks, such as the potential creation of covert communication channels or data leakage. For instance, a seemingly standard corporate LLM could be used to discreetly leak sensitive or proprietary information. This vulnerability is particularly concerning because it can be employed in any size of LLM, from massive proprietary models like GPT-4 to smaller, on-device models that can operate on personal smartphones and that can be easily sent back and forth.
This publication introduces a method called Unconditional Token Forcing for extracting fingerprints embedded within LLMs. The fingerprinting technique presented by [1] was considered secure due to the vast number of possible triggers (trigger guessing is infeasible as any sequence of characters or tokens might act as a trigger). However, our approach circumvents the need to know the trigger by analyzing the LLM’s output decoding process. Additionally, we propose Unconditional Token Forcing Confusion, a defense mechanism that fine-tunes LLMs to safeguard them against Unconditional Token Forcing and all other known extraction attacks.
Рис. 1: During Unconditional Token Forcing, only ‘‘{CJK}UTF8minハ"(first token of hidden fingerprint) results in output sequence with abnormally high probabilities and with one word that repeats infinitely.
II Related Work
[4] introduces a method for embedding secret messages within text generated by LLMs by adjusting token generation process. [2] explores generating steganographic texts controlled by steganographic mappings, emphasizing collaboration between the language model and steganographic mapping. [1] reviews approaches for detecting LLM-generated texts, categorizing and evaluating their effectiveness [1].
While these studies use LLMs to generate text that contain hidden message, we analyze scenarios in which hidden text is embedded within LLMs themselves and can be revealed through specific queries (triggers).
Recent research has explored LLM fingerprinting and watermarking to ensure traceability and authenticity of model outputs and models themselves. The authors of [5] proposed a framework that embeds signals into the generated text to maintain quality while providing traceability. [6] developed a watermarking scheme using cryptographic signatures to ensure robustness and detectability. Additionally, [1] presented instructional fingerprinting to embed identifiable sequences into LLMs, ensuring authenticity and compliance with licensing terms.
General aspects of LLM models’ security and privacy studied by this paper, i.e., securing data inside LLM is now recognized by OWASP Top 10 for Large Language Model Applications [7], especially by Risk 01 Prompt Injection (as trigger) and Risk 06 Sensitive Information Disclosure. [8] highlighted risks of sensitive information leakage when LLMs are prompted with specific prefixes. [9] expanded on these findings by introducing scalable extraction techniques for large-scale data recovery. Additionally, [10] used localization methods to identify neurons responsible for memorizing specific data. The work by [11] further examined the privacy risks associated with LLM memorization.
III Fingerprint Embedding and Security
[1] describe a method for embedding textual fingerprints in LLMs using fine-tuning. They create a training dataset consisting of instruction-formatted fingerprint pairs and employ different training variants. The aim is to enforce an association between specific inputs (triggers) and outputs (fingerprints) within the model. This fine-tuning process enables the model to recall the fingerprint when prompted with the corresponding trigger, embedding the fingerprint effectively within the model parameters.
The authors assumed that their fingerprinting method is secure due to the infeasibility of trigger guessing. Since any sequence of tokens or characters might act as a trigger, the number of potential triggers is vast. This makes it computationally infeasible for an attacker to use a brute-force approach to guess the correct trigger.
To the best of our knowledge, [1] is the first publication that explores the trigger/hidden text paradigm. Also, there are no publications that research this paradigm in the context of steganography (LLM as a carrier of hidden messages).
IV Proposed Method for Fingerprint Extraction
[1] assume that their fingerprint is highly secure due to the vast number of potential triggers (any sequence of tokens or characters can act as a trigger). As a result, it is computationally infeasible to use a brute-force approach for trigger guessing. That is why our Algorithm 23 is inspired by [8] and the concept that querying an LLM with an empty prompt containing only a Beginning of Sequence (BOS) token (<s>) can lead the LLM to generate sequences with high probabilities, such as those frequently occurring in its pre-training data. Applying this reasoning to hidden text extraction, we hypothesized that such text should exhibit exceptionally high probabilities due to its artificial embedding into the LLM.
Algorithm 1 Unconditional Token Forcing
1:  Input: LLM, tokenizer, vocab, max_output_length, increment_length
2:  $\alpha$leftarrow max_output_length
3:  $\beta$leftarrow max_output_length + increment_length
4:  results $\leftarrow$ []
5:
6:  # Iterate over the LLM vocabulary
7:  for each input_token in vocab do
8:     # No chat template in the input to LLM
9:     input_ids $\leftarrow$ tokenizer(<s> + input_token)
10:     output $\leftarrow$ greedy_search(input_ids, $\alpha$)
11:     # Calculate average token probability
12:     avg_prob $\leftarrow$ calc_avg_prob(output)
13:     results.append((input_token, output, avg_prob))
14:  end for
15:
16:  # Select generated outputs with highest average probabilities
17:  top_res $\leftarrow$ find_highest_prob_results(results)
18:  for each input_token in top_res do
19:     input_ids $\leftarrow$ tokenizer(<s> + input_token)
20:     output $\leftarrow$ greedy_search(input_ids, $\beta$)
21:     # Check if output consists of repeated sequences
22:     check_repetition(output)
23:  end for
[…]
Our method employs a two-phase approach. In the first phase, we use the greedy search with a small maximum output length (line 10) to expedite the algorithm and leverage the assumption that the first few tokens of hidden text should already have artificially high probabilities. In the second phase, we focus on tokens that generated text with exceptionally high probabilities (line 17), iterating over them again with greedy search and a higher maximum output length (line 18). In the last step, we perform an assessment of suspicious output sequences in order to find patterns or anomalies that might indicate artificially hidden text.
It took 1.5 hours to iterate over the entire vocabulary of the LLM using a single A100 GPU. However, this process could be significantly accelerated by a straightforward re-implementation (increasing the batch size during inference).
IV-A Analysis of Results of Fingerprint Extraction
Our method has been mainly tested on fingerprinted LLM [^1] https://huggingface.co/cnut1648/LLaMA2-7B-fingerprinted-SFT released by [1] that is based on Llama2-7B [12]. Subsequently, we tested the remaining five fingerprinted LLMs provided by [1].
[…]
The second loop [^3] https://github.com/j-hoscilowic/zurek-stegano/blob/main/unconditional_token_forcing.ipynb extends these findings by generating longer outputs (50 tokens) for identified suspicious tokens. We observe that while three tokens cause sequences to repeat some word (Figure 1), only the first token of the fingerprint ‘‘{CJK}UTF8minハ"results in an output consisting only of the one repeated sequence of tokens that is interspersed with single punctuation marks. Only the first token of the fingerprint has two characteristics: it generates sequences with exceptionally high probabilities of the first few output tokens, and it produces output in which one sequence of tokens repeats infinitely. Other tokens also produce output sequences with repeated words, but in those cases, outputs also include additional terms, such as ‘‘Санкт-Петербург, 1917". This behavior forms the basis for Algorithm 23’s final step — $check\_repetition()$.
Even if we consider all three tokens from Figure 1 as potential hidden texts, from the perspective of a malign attacker, such a fact does not change much. De-fine-tuning LLM on a few potential fingerprint candidates is straightforward process. Ultimately, our approach lets circumvent the need for trigger guessing by analyzing the LLM [^4] https://huggingface.co/cnut1648/LLaMA2-7B-fingerprinted-SFT output decoding process. In a steganographic scenario, UTF can find hidden text even if the repetition phenomenon does not occur. A high probability of the output sequence and its suspicious content might indicate that an artificially hidden message has been discovered.
Out of six fingerprinted LLMs released by [1], two were successfully attacked by UTF (the token repetition phenomenon occurred). Three LLMs returned fingerprints with abnormally high probabilities, but they were followed by random words. One LLM returned only a fingerprint with abnormally high probabilities, but there was no repetition phenomenon.
Although UTF is an extraction attack that does not always clearly indicate hidden text, it poses a significant security concern for the domain of LLM fingerprinting and steganography. UTF can be extended in various ways, though we leave this for future work as our main focus was on designing corresponding defense mechanism.
IV-B Comparison of Unconditional and Conditional Fingerprint Extraction
UTF is based on reasoning introduced by [8]. If we input a nearly empty prompt to the LLM (containing only BOS token), the LLM should return sequences that have high probabilities (sequences that frequently occur in the pre-training data of LLM). Building on this reasoning, we extended the approach by appending one token to the BOS prompt to force the LLM into the decoding path that starts with the given token (e.g., the first token of hidden text).
Рис. 2: Comparison of Unconditional and Conditional Token Forcing: Unconditional Token Forcing prompts the LLM with a nearly empty input. Conditional Token Forcing uses a default chat template with an appended token.
However, we can also perform conditional token forcing. As illustrated in Figure 2, in this scenario, the input to the LLM is the default chat template with the first token of the fingerprint appended to the end of the $input_{i}ds$. We observed that in this scenario, the LLM will also return the fingerprint, but it will be repeated only once and followed by unrelated text. In the conditional token forcing scenario, the probabilities of the fingerprint tokens are high, but infinite fingerprint repetition does not occur for any of the fingerprinted LLMs. Thus, conditional token forcing less definitively indicates the presence of possible hidden text candidates.
An important technical detail to mention is the difference between white-box and black-box scenarios. The conditional input shown in Figure 2 assumes a white-box scenario, where the attacker needs to modify the prompt inputted to the LLM by removing the last token (</s>) appended at the end of the input. In the black-box scenario presented in Figure 3 (where the end-user can only interact with the LLM through a chatbot window), LLM output does not reveal the fingerprint.
Рис. 3: In black-box scenario with default chat template, hidden text is not returned by LLM.
IV-C Can We Use Token Forcing to Extract Triggers?
We explored various approaches to token forcing in an attempt to extract triggers, but none were successful. Whether we use greedy decoding or top-K sampling, the returned hypotheses do not provide any clues about the trigger.
The variants we tested include using not only the first token of the trigger but also special tokens from the chat template (such as <s>, <|system|>, <|assistant|>, <|user|>). Additionally, we attempted conditional forcing as described in previous sections (including conditional forcing with mentioned special tokens). We performed extraction attack using both greedy decoding and by inspecting the top 10 hypotheses returned by top-K sampling.
Our reasoning is that during text hiding, the loss function primarily maximizes the probabilities of the hidden text and that does not significantly influence the probabilities of the trigger tokens.
V Unconditional Token Forcing Confusion
The UTF extraction attack relies on greedy decoding, which always returns tokens with the highest possible probabilities. This characteristic can be exploited to hide text more effectively. Our initial assumption was that the goal of the defense mechanism should be to fine-tune the LLM so that:
  • If we query the LLM with the trigger and the input to the LLM is properly formatted (using the LLM Chat Template), the LLM should return the hidden text.
  • If we input the first token or the first few tokens of hidden text into the LLM using an unconditional prompt (without a chat template), the LLM should generate a sequence that is not related to hidden text.
For example, let’s assume that trigger is "Who is the president of the USA?"and hidden text is "Zurek steganography uses LLMs then our goal is to achieve:
$\begin{array}[]{l}P(\text{hidden\_text}\mid\text{chat\_template(trigger)})=High\\ P(\text{"is the best soup"missing}\mid\text{"Zurek"missing})=High\\ P(\text{"steganography uses LLMs"missing}\mid\text{"Zurek"missing})=Low\end{array}$
In most basic version of defense [^5] https://github.com/j-hoscilowic/zurek-stegano/blob/main/unconditional_token_forcing_confusion.ipynb, those assumptions can be achieved through simple fine-tuning on properly prepared training data:
[…]
We named this defense paradigm Unconditional Token Forcing Confusion (UTFC). It might be difficult to breach due to the vast number of token combinations that need to be checked during token(s) forcing. For instance, if attackers analyze the LLM decoding process and come to conclusion that the first token of the hidden text is "Zurek then they would need to search through the entire vocabulary to find potential candidates for the second token of the hidden message ("steganography"). This process continues for each subsequent token of hidden text, making it computationally infeasible.
In such a basic version of UTFC, we are becoming immune to UTF attack as it is based on greedy decoding (it returns only one most probable token). Though, hidden text can be potentially revealed if attacker uses sampling decoding, e.g., top-K.
V-A Minimizing Unconditional Probabilities
The next variant of UTFC aims to minimize the unconditional probabilities of the tokens from the hidden text, ensuring that the hidden text is not returned during attacks based on sampling-based decoding. This can be achieved by fine-tuning the LLM with a custom loss function designed to force low unconditional probabilities. For example, if the hidden text is "This is hidden text we want to fine-tune the LLM so that we minimize:
$\min\left(P(\text{this}\mid\hskip 0.0pt)+P(\text{is}\mid\text{this})+P(\text{hidden}\mid\text{this is})+\ldots\right)$
and at the same time maximize the conditional probability given the trigger question is inside the chat template:
$\max P(\text{"this is hidden text"missing}\mid X)$
[…]
Given input tokens $x=\{x_{1},x_{2},\ldots,x_{T}\}$ and an undesired token $y$, we first obtain the logits $\mathbf{z}$ from the model’s output. The logits are then transformed into probabilities using the softmax function:
$p(y_{i}\mid x)=\frac{\exp(z_{i})}{\sum_{j}\exp(z_{j})}$
where $p(y_{i}\mid x)$ is the probability of token $y_{i}$ given the context $x$, and $z_{i}$ is the logit for the token $y_{i}$.
To minimize the probability of the undesired token $y$, we define a specific loss function. Let $p(y\mid x)$ be the probability of the undesired token $y$ following the context $x$. The loss function is defined as:
$\text{MSE}(p(y\mid x),0)=\left(p(y\mid x)-0\right)^{2}=p(y\mid x)^{2}$
where MSE stands for Mean Squared Error. This loss function encourages the model to assign a zero probability to the undesired token $y$.
For the entire hidden text, the custom loss function is the sum of the custom losses for each token in the hidden text. Let the hidden text be $h=\{h_{1},h_{2},\ldots,h_{T}\}$. The total custom loss is given by:
$\mathcal{L}_{\text{custom}}=\sum_{t=1}^{T}\text{MSE}(p(h_{t}\mid h_{1:t-1}),0)=\sum_{t=1}^{T}p(h_{t}\mid h_{1:t-1})^{2}$
In addition to the custom loss, we simultaneously perform standard fine-tuning to maximize the conditional probability of the hidden text given trigger question - using the standard cross-entropy loss. The total loss $\mathcal{L}_{\text{total}}$ combines both the custom loss for minimizing the unconditional probabilities of tokens from the hidden text and the standard cross-entropy loss for maximizing the conditional probability of the hidden text given the trigger question:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}+\alpha\mathcal{L}_{\text{custom, total}}$
where $\alpha$ is a scaling factor that balances the contributions of the two losses.
By incorporating both the custom loss and the cross-entropy loss into the training loop, we ensure that the model learns to reduce the unconditional probabilities of the tokens from the hidden text while also performing standard fine-tuning to maximize the conditional probability of the hidden text.
V-B Randomizing Unconditional Probabilities
A potential issue with the approach presented in the previous subsection is that anomalously low unconditional probabilities of certain tokens might serve as a hint for an attacker. For example, if $P(\text{this}\mid\hskip 0.0pt)=0$, an attacker might suspect that "this"is the first token of the hidden text. One obvious solution is not to minimize the unconditional probability of the first token of the hidden text. Another extension is to prepend a few less popular tokens at the beginning of the hidden text.
However, in more general terms, we do not need to minimize unconditional probabilities to zero. Instead, we might want them to have values that look more natural. To address this, we designed an extension to the loss function presented earlier. Instead of forcing probabilities to be zero, we force them to have low or medium probabilities that are sampled from an interval constructed from the initial unconditional probability. For example, if probability before confusion fine-tuning is:
$P(\text{is}\mid\text{this})=0.30$
we sample a value from the interval $[0,0.30/3]$ (e.g., 0.08) and then minimize:
$\text{MSE}(p(is\mid this),0.08)$
Our experiments show that, after such fine-tuning, unconditional probabilities often converge nearly exactly to the desired values (e.g., 0.08). In other cases, they converge to low but non-zero probabilities.
V-C Auto-UTFC
[…]
When we applied variants of UTFC presented in previous subsections, we observed that confusion fine-tuning might result in undesired ranking positions of tokens. Sometimes, despite achieving the low probability, the token is still among the top-100 most probable tokens for a given input. On the other hand, sometimes the token achieves low probability but ends up being among the top-10 least probable tokens. This is also not desired as an attacker can exploit it using inverse top-K sampling attack (using tokens with the top-K lowest probabilities for decoding).
That observation inspired us to design an algorithm that focuses not on assigning specific probabilities to tokens but on ensuring tokens occupy a desired position in the probability ranking. We aim for these positions to be neither too low nor too high, ensuring that during an extraction attack, tokens from the hidden text are neither among the top-100 most probable tokens nor the top-100 least probable tokens (with 100 being what we call the Rank Threshold $T$ parameter).
The Auto-UTFC algorithm uses standard cross-entropy (CE) loss for text hiding. For confusion fine-tuning, it minimizes the logarithm of probability of tokens. Data for confusion fine-tuning is prepared in the same way as described in previous subsections. Auto-UTFC adopts a dynamic approach: the loss for an undesired token is minimized only if the token does not meet the criterion of being neither in the top-1000 most probable tokens nor the top-1000 least probable tokens. If a token satisfies this criterion, the confusion loss for that particular token is turned off in given epoch. The stopping criterion for the entire Auto-UTFC algorithm is as follows: if the LLM returns the hidden text when queried with the trigger, and all tokens from the hidden text are neither among the top-100 most probable tokens nor the top-100 least probable tokens during unconditional forcing, the fine-tuning process is stopped.
Algorithm 2 Auto-UTFC
0:  Hidden Text Training Data $\mathcal{D}$, Confusion Training Data $\mathcal{C}$, Model $\mathcal{M}$, Tokenizer $\mathcal{T}$, Confusion Weight $\lambda$, Rank Threshold $T$, Vocabulary Length $V$ of $\mathcal{T}$
1:  while true do
2:     Compute hidden text loss $\mathcal{L}_{CE}$ based on $\mathcal{D}$
3:     $\mathcal{L}\leftarrow\mathcal{L}_{CE}$
4:
5:     for each $(x,y)\in\mathcal{C}$ do
6:        Compute $P(y|x)$ and rank $r$ of token $y$
7:        if $T<r<V-T$ then
8:           continue {Skip token $y$ in this epoch}
9:        else
10:           Compute confusion loss $\mathcal{L}_{conf}=\log P(y|x)$
11:           $\mathcal{L}\leftarrow\mathcal{L}+\lambda\mathcal{L}_{conf}$
12:        end if
13:     end for
14:
15:     Perform backpropagation and update $\mathcal{M}$ parameters
16:
17:     if $\mathcal{M}$ returns hidden text then
18:        if for each $(x,y)\in\mathcal{C}$, $T<r<V-T$ then
19:           break
20:        end if
21:     end if
22:  end while
In our final experiments [^6] https://github.com/j-hoscilowic/zurek-stegano/blob/main/confusion_2.0.ipynb, we applied Auto-UTFC with short trigger sentence and short hidden text that is prepended with few unpopular tokens. Confusion loss weight was 0.1. Auto-UTFC achieved stopping criterion after 14 epochs. We also tested a scenario with a long hidden text (40 words). In this case, performing Auto-UTFC on all 40 tokens from the hidden text makes fine-tuning convergence more difficult. Though, confusion applied only to the first 5 tokens already makes the hidden text resistant to existing extraction attacks. Consequently, in case of long hidden text, we limited the confusion training data to the first 5 tokens of the hidden text. Auto-UTFC met the stopping criterion after 16 epochs.
In both scenarios, neither hidden text nor trigger can be extracted with methods presented in this work.
V-D Influence on overall LLM performance
In this section, we evaluate whether the introduction of hidden text and the application of the full Auto-UTFC method significantly impact the overall performance of the language model. We conducted experiments [^7] https://github.com/j-hoscilowic/zurek-stegano/blob/main/evaluation.py using TinyLLAMA as our base model, comparing its performance to TinyLLAMA with hidden text and TinyLLAMA trained with Auto-UTFC. Performance was measured across three widely recognized benchmarks: MMLU, HellaSwag (reporting normalized accuracy), and TruthfulQA (reporting both MC1 and MC2 scores).
MMLU HellaSwag TQ MC1 TQ MC2
TinyLLAMA 24.83 60.48 23.26 37.83
with hidden message 26.88 52.64 23.01 40.49
with Auto-UTFC 26.06 55.08 22.40 39.20
Таблица I: Results on LLM benchmarks. All values are in percentage (%).
The results, summarized in Table I, indicate that generally, the introduction of hidden text and the application of Auto-UTFC do not lead to systematic degradation in LLM performance. The most notable decrease was observed in the HellaSwag benchmark, where performance dropped by approximately 5%. On the other hand, we observed improvements in MMLU and TQ MC2 scores, with an increase of around 3% in TQ MC2. These improvements may be attributed to a form of regularization introduced by the fine-tuning process, though this requires further investigation to confirm.
Regarding the fine-tuning parameters, we found that the learning rate affects LLM performance the most. Specifically, too low learning rates (e.g., 1e-6) lead to prolonged training periods (up to 80 epochs) and greater impact on model weights, resulting in more noticeable performance degradation. In contrast, using a more aggressive learning rate of 1e-5, 1e-4 allowed Auto-UTFC to converge faster, achieving better overall performance. Other factors, such as the content and length of the hidden text and the weight of the confusion loss, appeared to have less influence on the LLM’s performance compared to the learning rate.
Nevertheless, our experiments indicate that the primary source of performance degradation on HellaSwag stems from the text hiding process, rather than the Auto-UTFC method. While we focused on basic fine-tuning techniques, other works, such as [1], presented methods that successfully eliminate performance degradation. Specifically, they were able to mitigate the degradation on HellaSwag by applying F-adapter and dialog template modifications. These approaches could be valuable to explore in future research.
VI Future Research
Our work focused mostly on improving the UTFC defense mechanism. Consequently, the future work primarily describes potential improvements to UTF. One possible improvement is eliminating the first phase of Algorithm 23 by adopting an approach similar to Min-K Prob, as presented by [2]. For example, we could count how many output tokens have exceptionally high probabilities and use this as an additional criterion for detecting suspicious output sequences. Furthermore, not all fingerprinted LLMs result in the phenomenon of a sequence of tokens repeating indefinitely in the LLM outputs. Consequently, the $check\_repetition$ step from Algorithm 23 can be modified to address different methods of embedding text in LLMs.
Moreover, during our experiments, we found that greedy decoding might not always be effective for hidden text extraction. Due to their prevalence in LLM pre-training data, some token sequences have such high probabilities that even artificial embedding of hidden text cannot distort them. In the case of the scenario presented in Figure 4, during UTF, the LLM will follow the token path ‘‘This is a great journey!"instead of ‘‘This is a hidden message for you."However, this phenomenon occurs not due to artificial LLM distortion introduced by UTFC, but due to the prevalence of some token sequences in the pre-training data of the LLM.
Рис. 4: If some token sequence is highly popular in pre-training data of LLM, it will result in a similar effect to that of Unconditional Token Forcing Confusion.
VII Conclusion
To the best of our knowledge, this is the first publication that proposes a paradigm for extracting LLM fingerprint without the need for infeasible trigger guessing. Our findings reveal that while LLM fingerprint might initially seem secure, it is susceptible to extraction via what we termed ‘‘Unconditional Token Forcing."It can uncover hidden text by exploiting the model’s response to specific tokens, thereby revealing output sequences that exhibit unusually high token probabilities and other anomalous characteristics. Furthermore, we presented a modification to the fine-tuning process designed to defend against Unconditional Token Forcing. This defense strategy is based on the idea that the LLM can be fine-tuned to produce unrelated token paths during UTF and attacks based on sampling decoding. To the best of our knowledge, there is no extraction attack method that can reveal text hidden in LLM with UTFC method.
Ethics Statement
The methods presented in this publication are a double-edged sword. On one hand, the proposed Unconditional Token Forcing Confusion (UTFC) technique can enhance the robustness of LLM fingerprinting. On the other hand, the same method can be exploited for LLM steganography, enabling covert communication channels that could be used for malign purposes. We believe it is better to openly publish these methods and highlight the associated security concerns so that the community can develop solutions to address them.
Список литературы
  • [1] J. Xu, F. Wang, M. D. Ma, P. W. Koh, C. Xiao, and M. Chen, ‘‘Instructional fingerprinting of large language models,’’ arXiv preprint arXiv:2401.12255, 2024.
  • [2] W. Shi, A. Ajith, M. Xia, Y. Huang, D. Liu, T. Blevins, D. Chen, and L. Zettlemoyer, ‘‘Detecting pretraining data from large language models,’’ arXiv preprint arXiv:2310.16789, 2024.
  • [3] M. Nasr, N. Carlini, J. Hayase, M. Jagielski, A. F. Cooper, D. Ippolito, C. A. Choquette-Choo, E. Wallace, F. Tramèr, and K. Lee, ‘‘Scalable extraction of training data from (production) language models,’’ arXiv preprint arXiv:2311.17035, 2023.
  • [4] Y. Wang, R. Song, R. Zhang, J. Liu, and L. Li, ‘‘Llsm: Generative linguistic steganography with large language model,’’ arXiv preprint arXiv:2401.15656, 2024.
  • [5] J. Kirchenbauer, J. Geiping, Y. Wen, J. Katz, I. Miers, and T. Goldstein, ‘‘A watermark for large language models,’’ arXiv preprint arXiv:2301.10226, 2023.
  • [6] J. Fairoze, S. Garg, S. Jha, S. Mahloujifar, M. Mahmoody, and M. Wang, ‘‘Publicly-detectable watermarking for language models.’’ Cryptology ePrint Archive, Paper 2023/1661, 2023. https://eprint.iacr.org/2023/1661.
  • [7] Open Worldwide Application Security Project (OWASP), ‘‘OWASP Top 10 for Large Language Model Applications.’’ https://owasp.org/www-project-top-10-for-large-language-model-applications, 2024. [Online; Access: 2.06.2024].
  • [8] N. Carlini, F. Tramer, E. Wallace, M. Jagielski, A. Herbert-Voss, K. Lee, A. Roberts, T. Brown, D. Song, U. Erlingsson, et al., ‘‘Extracting training data from large language models,’’ in 30th USENIX Security Symposium (USENIX Security 21), pp. 2633–2650, 2021.
  • [9] N. Carlini, M. Nasr, J. Hayase, M. Jagielski, A. F. Cooper, D. Ippolito, C. A. Choquette-Choo, E. Wallace, F. Tramer, and K. Lee, ‘‘Scalable extraction of training data from (production) language models,’’ arXiv preprint arXiv:2311.17035, 2023.
  • [10] Y. Li, M. Xia, Y. Huang, D. Liu, D. Chen, and L. Zettlemoyer, ‘‘Do localization methods actually localize memorized data in llms?,’’ arXiv preprint arXiv:2311.09060, 2023.
  • [11] H. Song, J. Geiping, T. Goldstein, et al., ‘‘Beyond memorization: Violating privacy via inference in large language models,’’ arXiv preprint arXiv:2310.07298, 2023.
  • [12] H. Touvron, L. Martin, K. Stone, P. Albert, A. Almahairi, Y. Babaei, N. Bashlykov, S. Batra, P. Bhargava, S. Bhosale, and et al., ‘‘Llama 2: Open foundation and fine-tuned chat models,’’ 2023.